# Imbalance-Robust and Sampling-Efficient Continuous Conditional GANs via Adaptive Vicinity and Auxiliary Regularization

### v2

@@ Abstract
Recent advances in conditional generative modeling have introduced Continuous conditional Generative Adversarial Network (CcGAN) and Continuous Conditional Diffusion Model (CCDM) for estimating high-dimensional data distributions conditioned on scalar, continuous regression labels (e.g., angles, ages, or temperatures). However, these approaches face fundamental limitations: CcGAN suffers from data imbalance due to fixed-size vicinity constraints, while CCDM requires computationally expensive iterative sampling. We present CcGAN-AVAR, an enhanced CcGAN framework that addresses both challenges: (1) leveraging the GAN framework’s native one-step generation to overcome CCDMs’ sampling bottleneck (achieving 300×-2000× faster inference), while (2) two novel components specifically target data imbalance - an adaptive vicinity mechanism that dynamically adjusts vicinity’s size, and a multi-task discriminator that constructs two regularization terms (through auxiliary regression and density ratio estimation) to significantly improve generator training. Extensive experiments on four benchmark datasets (64×64 to 192×192 resolution) across eight challenging imbalanced settings demonstrate that CcGAN-AVAR achieves state-of-the-art generation quality while maintaining sampling efficiency.
Codehttps://github.com/UBCDingXin/CcGAN-AVAR
1 Introduction
@@
Figure 1: The illustration of CCGM with data imbalance in UTKFace and Steering Angle datasets.
@@
To mitigate data scarcity, Ding et al. (2021, 2023b) proposed the Continuous conditional Generative Adversarial Network (CcGAN), which estimates the target data distribution $p_{r}(\bm{x}|y_{c})$ by incorporating training samples within a fixed-size hard or soft vicinity of the conditional label $y_{c}$. While these vicinities enable CcGAN to trade some label consistency—the degree to which generated samples align with the given condition—for improved training stability and generation quality, Figure 2(a)–(b) reveals a critical limitation: the fixed-size vicinities may unnecessarily compromise too much label consistency for densely sampled label values, ultimately leading to suboptimal generation performance.
@@
As an alternative to CcGAN, Ding et al. (2025) recently integrated CcGAN techniques into diffusion models, proposing the Continuous Conditional Diffusion Model (CCDM). Although CCDM exhibits superior generation quality and robustness to data imbalance compared to CcGAN, its sampling process is 700–2000× slower, severely limiting practical deployment. While distillation-based one-step sampling mitigates this latency, it incurs significant generation quality degradation and additional training overhead (Ding et al. 2025). Thus, achieving both high-quality generation and efficient sampling remains an open challenge in CCGM.
Motivated by these limitations, we present CcGAN-AVAR, an improved CcGAN framework that achieves generation quality comparable to or surpassing CCDM while maintaining significantly faster sampling speeds (300×-2000× acceleration). Our key contributions include: (1) a novel soft/hybrid Adaptive Vicinity (AV) mechanism that overcomes the limitations of conventional hard/soft vicinities in handling data imbalance; (2) a multi-task discriminator with dedicated auxiliary branches for noise-insensitive regression and density ratio estimation, generating two auxiliary regularization terms that substantially improve generation quality and label consistency; (3) RC-49-I, an imbalanced variant of the RC-49 dataset (Ding et al. 2023b) incorporating three distinct imbalance patterns (unimodal, bimodal, and trimodal distributions) for thorough evaluation of CCGM methods under realistic data scarcity conditions; and (4) an optimized, unified codebase for CcGAN supporting multiple benchmark datasets, flexible network architecture and vicinity selection, various loss functions, exponential moving average, and mixed-precision training.
2 Related Work
2.1 Continuous Conditional Generative Models
@@
To address data scarcity in CCGM tasks, CcGAN (Ding et al. 2021, 2023b) leverages samples within a label-dependent vicinity of $y_{c}$ to approximate the data distribution $p_{r}(\bm{x}|y_{c})$. The framework introduces two distinct vicinity formulations (hard vicinity and soft vicinity), which leads to the Hard/Soft Vicinal Discriminator Loss (HVDL/SVDL) and the corresponding generator loss for CcGAN:
$\mathcal{L}(D)=$
@@
$-\frac{1}{N^{g}}\sum_{j=1}^{N^{g}}\sum_{i=1}^{N^{g}}\mathds{E}_{\epsilon\sim\mathcal{N}(0,\sigma^{2})}\left[W^{g}_{i}\log(1-D(\bm{x}_{i}^{g},y_{j}^{g}+\epsilon))\right],$ (1)
$\mathcal{L}(G)=$
@@
where $D$ and $G$ denote the discriminator and generator networks, superscripts rr/$g$ indicate real/fake samples, $\bm{z}_{i}$ represents Gaussian noise, $\sigma$ is a positive hyperparameter, and $N^{r}$$N^{g}$ denote real/fake sample sizes. As illustrated in Figure 2(a), the hard vicinity defines a closed interval centered at $y_{c}$ with a fixed radius $\kappa$. Thus, the weights $W^{r}_{i}$ and $W^{g}_{i}$ for HVDL are defined as follows:
@@
with $\mathds{1}_{\{\cdot\}}$ as the indicator function and $N_{y_{c},\kappa}$ counting samples in the hard vicinity $y_{c}-\kappa,y_{c}+\kappa]$. Ding et al. (2023b) provides a rule of thumb to compute $\kappa$ and $\sigma$ from samples. For soft vicinity, all samples are used to estimate $p_{r}(\bm{x}|y_{c})$, but each sample $\bm{x}_{i}$ is assigned with a weight $W_{i}$, with the following formulation:
@@
As illustrated in Figure 2(b), the soft vicinity weights in Eq. (4) follow an exponential decay curve $w(y,y^{\prime})=e^{-\nu(y-y^{\prime})^{2}}$, where the decay rate $\nu>0$ controls the weighting sensitivity to label distances. Ding et al. (2023b) set $\nu=1/\kappa^{2}$ and exclude samples with soft weights below $10^{-3}$ from discriminator updates for computational efficiency.
@@
The conventional hard/soft vicinity (with constant $\kappa$ and $\nu$) often produces suboptimal generation quality on imbalanced datasets. As Figure 2(a)-(b) shows, while samples cluster densely around $y_{c}^{2}$, the fixed-size vicinity formulation either incorporates excessive samples (hard vicinity) or over-weights distant samples (soft vicinity), both of which significantly degrade label consistency.
Recent work by Ding et al. (2025) introduces CCDM, which incorporates vicinal techniques into diffusion models as a replacement for CcGAN. Although both their empirical study and the experiments in this paper (Tables 1 and 2) show CCDM outperforms conventional CcGANs while maintaining data imbalance robustness, its prohibitively slow sampling speed significantly limits practical applicability.
2.2 Sample-Based Density Ratio Estimation
Ding et al. (2023a) introduced cDR-RS, a sample-based Density Ratio Estimation (DRE) method for conditional GANs that computes $r(\bm{x}|y)=p_{r}(\bm{x}|y)/p_{g}(\bm{x}|y)$. The estimated density ratios enable rejection sampling to enhance output quality in both class-conditional GANs and CcGANs.
cDR-RS’s DRE framework employs a two-stage architecture: a fixed, pre-trained encoder that maps input images $\bm{x}$ into latent representations $\bm{h}$, and a trainable 5-layer fully-connected network $f_{\text{dre}}$ that processes $\bm{h}$ for density ratio estimation. During optimization, only $f_{\text{dre}}$ is updated using the following penalized Softplus loss:
$\mathcal{L}(f_{\text{dre}})=\frac{1}{N^{g}}\sum_{i=1}^{N^{g}}\left[\delta(f_{\text{dre}}(\bm{h}_{i}^{g}|y_{i}^{g}))f_{\text{dre}}(\bm{h}_{i}^{g}|y_{i}^{g})-\eta(f_{\text{dre}}(\bm{h}_{i}^{g}|y_{i}^{g}))\right]$
$-\frac{1}{N^{r}}\sum_{i=1}^{N_{r}}\delta(f_{\text{dre}}(\bm{h}_{i}^{r}|y_{i}^{r}))+\lambda_{\text{dre}}\cdot\left(\frac{1}{N^{g}}\sum_{i=1}^{N^{g}}f_{\text{dre}}(\bm{h}_{i}^{g}|y_{i}^{g})-1\right)^{2},$ (5)
@@ 2.2 Sample-Based Density Ratio Estimation
3 Method
3.1 Overview
In this section, we present CcGAN-AVAR, an improved CcGAN framework designed to address data imbalance challenges. As illustrated in Figure 3, our approach introduces two key innovations: (1) a novel soft/hybrid adaptive vicinty mechanism that dynamically adjusts to local sample density, and (2) a multi-task discriminator that generates two auxiliary regularization terms to enhance generator training.
Figure 3: Overall workflow of CcGAN-AVAR
3.2 Adaptive Vicinity for Imbalanced Data
Let $Q^{r}=\{(\bm{x}^{r}_{i},y^{r}_{i})\}_{i=1}^{N^{r}}$ denote a training set of $N^{r}$ real image-label pairs, where the labels consist of $M$ distinct values $Y^{u}=\{y_{(1)},...,y_{(M)}\}$ sorted in strictly increasing order ($y_{(i)}<y_{(i+1)}$ for i=1,,Mi=1,...,M). For $i$-th distinct label value $y_{(i)}$, we denote by $N_{i}$ its corresponding sample count. In severely imbalanced datasets, $N_{i}$ exhibits substantial variation across different $i$, as demonstrated in Figure 1(b)-(c).
To address data imbalance, we propose two distinct types of Adaptive Vicinity (AV): Soft AV and Hybrid AV. Both approaches replace the vicinity parameters $\kappa$ and $\nu$ in Eqs. (3) and (4) with a new hyperparameter $N_{\text{AV}}$, which determines the minimum number of effective samples per vicinity. Practical guidance for setting $N_{\text{AV}}$ is provided in the Appendix.
@@ 3.2 Adaptive Vicinity for Imbalanced Data
Building on CcGAN’s empirically demonstrated superiority of soft vicinity over hard vicinity (Ding et al. 2021, 2023b), we first propose a soft AV mechanism (Figure 2(c)). During discriminator training, we create a target conditional label $y_{c}$ by randomly adding Gaussian noise to a training label (e.g., $y_{j}^{r}+\epsilon$ in Eq. (1)). Without loss of generality, assuming $y_{c}$ lies between $y_{(i-1)}$ and $y_{(i)}$, we initialize a hard vicinity centered at $y_{c}$ with zero-radius boundaries ($\kappa_{l}=0$, $\kappa_{r}=0$) containing $N_{c}=0$ samples. This vicinity then extends iteratively until $N_{c}\geq N_{\text{AV}}$: at each step, the vicinity extends leftwards if the leftmost external label is closer to $y_{c}$ than the rightmost external label, and rightwards otherwise. Following each extension, we update either $\kappa_{l}$ or $\kappa_{r}$ to reflect the distance of the newly incorporated label to $y_{c}$, while incrementing $N_{c}$ by the sample count at this label. Upon reaching $N_{c}\geq N_{\text{AV}}$, we obtain the final radius $\kappa_{y_{c}}=\max(\kappa_{l},\kappa_{r})$, which determines the local weight decay rate $\nu_{y_{c}}=1/\kappa_{y_{c}}^{2}$. Soft weights are then assigned to all training samples via Eq. (4) using $\nu_{y_{c}}$, with Figure 4 illustrating this process.
This adaptive approach automatically adjusts to local sample density: in dense regions (small $\kappa_{y_{c}}$), weights decay rapidly to maintain label consistency while preserving diversity; in sparse regions (large $\kappa_{y_{c}}$), weights decay slowly to incorporate more distant samples, ensuring training stability and preventing mode collapse.
@@
We further propose a hybrid AV mechanism (Figs. 2(d) and 4) that extends the soft AV approach. This hybrid variant incorporates truncated weight decay by assigning zero weights to samples outside the interval $y_{c}-\kappa_{y_{c}},y_{c}+\kappa_{y_{c}}]$. By combining the adaptive weighting of soft AV with the boundary constraints of hard vicinity, this approach can further enhances label consistency.
3.3 Auxiliary Generator Regularization

### v3

@@ Abstract
Recent advances in conditional generative modeling have introduced Continuous conditional Generative Adversarial Network (CcGAN) and Continuous Conditional Diffusion Model (CCDM) for estimating high-dimensional data distributions conditioned on scalar, continuous regression labels (e.g., angles, ages, or temperatures). However, these approaches face fundamental limitations: CcGAN suffers from data imbalance due to fixed-size vicinity constraints, while CCDM requires computationally expensive iterative sampling. To address these issues, we propose CcGAN-AVAR, an enhanced CcGAN framework featuring (1) two novel components for handling data imbalance—an adaptive vicinity mechanism that dynamically adjusts vicinity size and a multi-task discriminator that enhances generator training through auxiliary regression and density ratio estimation—and (2) the GAN framework’s native one-step generation, enabling 300×–2000× faster inference than CCDM. Extensive experiments on four benchmark datasets (64×64 to 256x256 resolution) across eleven challenging settings demonstrate that CcGAN-AVAR achieves state-of-the-art generation quality while maintaining sampling efficiency. Our code will be made publicly available at https://github.com/UBCDingXin/CcGAN-AVAR.
I Introduction
@@
Figure 1: The illustration of CCGM with data imbalance in UTKFace and Steering Angle datasets. Fig. 1(a) is adapted from Fig. 1 in [14].
@@
To mitigate data scarcity, Ding et al. [15, 16, 17] proposed the Continuous conditional Generative Adversarial Network (CcGAN), which estimates the target data distribution $p_{r}(\bm{x}|y_{c})$ by incorporating training samples within a fixed-size hard or soft vicinity of the conditional label $y_{c}$. While these vicinities enable CcGAN to trade some label consistency—the degree to which generated samples align with the given condition—for improved training stability and generation quality, Fig. 2(a)–(b) reveals a critical limitation: the fixed-size vicinities may unnecessarily compromise too much label consistency for densely sampled label values, ultimately leading to suboptimal generation performance.
@@
As an alternative to CcGAN, Ding et al. [14] recently integrated CcGAN techniques into diffusion models, proposing the Continuous Conditional Diffusion Model (CCDM). Although CCDM exhibits superior generation quality and robustness to data imbalance compared to CcGAN, its sampling process is 700–2000× slower, severely limiting practical deployment. While distillation-based one-step sampling mitigates this latency, it incurs significant generation quality degradation and additional training overhead [14]. Thus, achieving both high-quality generation and efficient sampling remains an open challenge in CCGM.
Motivated by these limitations, we present CcGAN-AVAR, an improved CcGAN framework that achieves generation quality comparable to or surpassing CCDM while maintaining significantly faster sampling speeds (300×-2000× acceleration). Our key contributions include:
  • A novel Soft/Hybrid Adaptive Vicinity (SAV/HAV) mechanism that overcomes the limitations of conventional hard/soft vicinities in handling data imbalance.
  • A multi-task discriminator with dedicated auxiliary branches for noise-insensitive regression and density ratio estimation, introducing two auxiliary regularization terms during generator training that substantially improve generation quality and label consistency.
  • An unified, scalable codebase for CcGAN and CcGAN-AVAR supporting multiple benchmark datasets, flexible network architectures and vicinity selection, exponential moving average, and mixed-precision training.
  • RC-49-I, an imbalanced variant of the RC-49 dataset [16], incorporating three distinct imbalance patterns (unimodal, bimodal, and trimodal distributions) to enable comprehensive evaluation of continuous conditional generative models under realistic data scarcity conditions.
  • The $256\times 256$ versions of the UTKFace and Steering Angle datasets for validating the effectiveness of CcGAN-AVAR in more challenging high-resolution scenarios.
II Preliminary
II-A Continuous Conditional Generative Models
@@
To address data scarcity in CCGM tasks, CcGAN [15, 16] leverages samples within a label-dependent vicinity of $y_{c}$ to approximate the data distribution $p_{r}(\bm{x}|y_{c})$. The framework introduces two distinct vicinity formulations (hard vicinity and soft vicinity), which leads to the Hard/Soft Vicinal Discriminator Loss (HVDL/SVDL) and the corresponding generator loss for CcGAN:
$\mathcal{L}^{D}_{\text{adv}}=$
@@
$-\frac{1}{N^{g}}\sum_{j=1}^{N^{g}}\sum_{i=1}^{N^{g}}\mathds{E}_{\epsilon\sim\mathcal{N}(0,\sigma^{2})}\left[W^{g}_{i}\log(1-D(\bm{x}_{i}^{g},y_{j}^{g}+\epsilon))\right]$ (1)
$\mathcal{L}^{G}_{\text{adv}}=$
@@
where $D$ and $G$ denote the discriminator and generator networks, superscripts rr/$g$ indicate real/fake samples, $W^{r}_{i}$ and $W^{g}_{i}$ are the weights for the training samples, $\bm{z}_{i}$ represents Gaussian noise, $\sigma$ is a positive hyperparameter, and $N^{r}$$N^{g}$ denote real/fake sample sizes. These CcGAN losses are derived from the vanilla GAN objective [18]; their hinge loss variants are provided in Eq. (S.26) of the Appendix in [16].
As illustrated in Fig. 2(a), the hard vicinity defines a closed interval centered at $y_{c}$ with a fixed radius $\kappa$. Thus, the weights $W^{r}_{i}$ and $W^{g}_{i}$ for HVDL are defined as follows:
@@
with $\mathds{1}_{\{\cdot\}}$ as the indicator function and $N_{y_{c},\kappa}$ counting samples in the hard vicinity $y_{c}-\kappa,y_{c}+\kappa]$. Ding et al. [16] provide a rule of thumb to compute $\kappa$ and $\sigma$ from samples. For soft vicinity, all samples are used to estimate $p_{r}(\bm{x}|y_{c})$, but each sample $\bm{x}_{i}$ is assigned with a weight $W_{i}$, with the following formulation:
@@
As illustrated in Fig. 2(b), the soft vicinity weights in Eq. (4) follow an exponential decay curve
$w(y,y^{\prime})=e^{-\nu(y-y^{\prime})^{2}}$ (5)
where the decay rate $\nu>0$ controls the weighting sensitivity to label distances. Ding et al. [16] set $\nu=1/\kappa^{2}$ and exclude samples with soft weights below $10^{-3}$ from discriminator updates for computational efficiency.
@@
The conventional hard/soft vicinity (with constant $\kappa$ and $\nu$) often produces suboptimal generation quality on imbalanced datasets. As Fig. 2(a)-(b) shows, while samples cluster densely around $y_{c}^{2}$, the fixed-size vicinity formulation either incorporates excessive samples (hard vicinity) or over-weights distant samples (soft vicinity), both of which significantly degrade label consistency.
Besides the vicinal techniques, to address the challenges of integrating continuous regression labels into conditional generative models, Ding et al. introduced an Improved Label Input (ILI) mechanism, which transforms an one-dimensional regression label $y$ into a multi-dimensional embedding vector $\bm{h}_{y}=\phi(y)$ via a pre-trained 5-layer multi-layer perceptron (MLP) $\phi$, enabling effective injection of the encoded label information into both the generator and discriminator networks.
Recent work by [14] introduces CCDM, which incorporates vicinal techniques and ILI’s embedding method into diffusion models as a replacement for CcGAN. Although both their empirical study and the experiments in this paper (Tables I and II) show CCDM outperforms conventional CcGANs while maintaining data imbalance robustness, its prohibitively slow sampling speed significantly limits practical applicability. To address this, [14] adapted a distillation-based technique [19], proposing the one-step DMD2-M sampler. Although DMD2-M achieves significantly faster inference than CCDM’s default DDIM sampler [20], it exhibits non-negligible quality degradation, as demonstrated in [14]’s experiments.
Thus, developing a sampling-efficient CCGM method that maintains imbalance robustness remains an open challenge.
II-B Sample-Based Density Ratio Estimation
Ding et al. [21] introduced cDR-RS, a sample-based Density Ratio Estimation (DRE) method for conditional GANs that computes $r(\bm{x}|y)=p_{r}(\bm{x}|y)/p_{g}(\bm{x}|y)$. The estimated density ratios enable rejection sampling to enhance output quality in both class-conditional GANs and CcGANs.
cDR-RS’s DRE framework employs a two-stage architecture: a fixed, pre-trained encoder that maps input images $\bm{x}$ into latent representations $\bm{h}$, and a trainable 5-layer linear network $f_{\text{dre}}$ that processes $\bm{h}$ for density ratio estimation. During optimization, only $f_{\text{dre}}$ is updated using the following penalized Softplus loss [22]:
$\mathcal{L}(f_{\text{dre}})=$
$\frac{1}{N^{g}}\sum_{i=1}^{N^{g}}\left[\delta(f_{\text{dre}}(\bm{h}_{i}^{g}|y_{i}^{g}))f_{\text{dre}}(\bm{h}_{i}^{g}|y_{i}^{g})-\eta(f_{\text{dre}}(\bm{h}_{i}^{g}|y_{i}^{g}))\right]$
$-\frac{1}{N^{r}}\sum_{i=1}^{N_{r}}\delta(f_{\text{dre}}(\bm{h}_{i}^{r}|y_{i}^{r}))$
$+\lambda_{\text{dre}}\cdot\left(\frac{1}{N^{g}}\sum_{i=1}^{N^{g}}f_{\text{dre}}(\bm{h}_{i}^{g}|y_{i}^{g})-1\right)^{2}$ (6)
@@ II-B Sample-Based Density Ratio Estimation
III Method
III-A Overview
In this section, we present CcGAN-AVAR, an improved CcGAN framework designed to address data imbalance challenges and serve as an efficient alternative to CCDM by enabling significantly faster inference. As illustrated in Fig. 3, our approach introduces two key innovations: (1) a novel soft/hybrid adaptive vicinty mechanism (introduced in Section III-B) that dynamically adjusts to local sample density, and (2) a multi-task discriminator (described in Section III-C) that generates two auxiliary regularization terms to enhance generator training. The neural network architectures are detailed in Section III-D, and a unified, scalable codebase for CcGAN-AVAR—which differs from prior CcGAN implementations [15, 16, 23, 17, 21]—is described in Section III-E.
Figure 3: Training pipeline of the CcGAN-AVAR framework. Novel components and procedures are highlighted in red. Data flow and loss computations are depicted by solid and dashed lines, respectively. The regression label $y_{c}$ is converted into an embedding vector $\bm{h}_{y_{c}}$ via the Improved Label Input (ILI) mechanism [16]. Trainable and frozen modules are distinctly marked by flame and frost icons.
III-B Adaptive Vicinity for Imbalanced Data
Let $Q^{r}=\{(\bm{x}^{r}_{i},y^{r}_{i})\}_{i=1}^{N^{r}}$ denote a training set of $N^{r}$ real image-label pairs, where the labels consist of $M$ distinct values $Y^{u}=\{y_{(1)},...,y_{(M)}\}$ sorted in strictly increasing order ($y_{(i)}<y_{(i+1)}$ for i=1,,Mi=1,...,M). For $i$-th distinct label value $y_{(i)}$, we denote by $N_{i}$ its corresponding sample count. In severely imbalanced datasets, $N_{i}$ exhibits substantial variation across different $i$, as demonstrated in Fig. 1(b)-(c).
To address data imbalance, we propose two distinct types of Adaptive Vicinity (AV): Soft AV (SAV) and Hybrid AV (HAV). Both approaches replace the vicinity parameters $\kappa$ and $\nu$ in Eqs. (3) and (4) with a new hyperparameter $N_{\text{AV}}$, which determines the minimum number of effective samples per vicinity. Practical guidance for setting $N_{\text{AV}}$ is provided in the Remark 2.
@@ III-B Adaptive Vicinity for Imbalanced Data
Building on CcGAN’s empirically demonstrated superiority of soft vicinity over hard vicinity [15, 16], we first propose a soft AV mechanism (Fig. 2(c)). During discriminator training, we create a target conditional label $y_{c}$ by randomly adding Gaussian noise to a training label (e.g., $y_{j}^{r}+\epsilon$ in Eq. (1)). Without loss of generality, assuming $y_{c}$ lies between $y_{(i-1)}$ and $y_{(i)}$, we initialize a hard vicinity centered at $y_{c}$ with zero-radius boundaries ($\kappa_{l}=0$, $\kappa_{r}=0$) containing $N_{c}=0$ samples. This vicinity then extends iteratively until $N_{c}\geq N_{\text{AV}}$: at each step, the vicinity extends leftwards if the leftmost external label is closer to $y_{c}$ than the rightmost external label, and rightwards otherwise. Following each extension, we update either $\kappa_{l}$ or $\kappa_{r}$ to reflect the distance of the newly incorporated label to $y_{c}$, while incrementing $N_{c}$ by the sample count at this label. Upon reaching $N_{c}\geq N_{\text{AV}}$, the final radius $\kappa_{y_{c}}=\max(\kappa_{l},\kappa_{r})$ determines the local weight decay rate $\nu_{y_{c}}=1/\kappa_{y_{c}}^{2}$, which yields the local exponential decay curve:
$w_{y_{c}}(y,y_{c})=e^{-\nu_{y_{c}}(y-y_{c})^{2}}.$ (7)
Substituting Eq. (7) into Eq. (4), the soft adaptive weight for the $i$-th real/fake training sample is computed as:
$W_{i,y_{c}}=\frac{w_{y_{c}}(y_{i},y_{c})}{\sum_{i=1}^{N}w_{y_{c}}(y_{i},y_{c})}=\frac{e^{-\nu_{y_{c}}(y_{i}-y_{c})^{2}}}{\sum_{i=1}^{N}e^{-\nu_{y_{c}}(y_{i}-y_{c})^{2}}},$ (8)
where $W_{i,y_{c}}$ depends on both the distance of $y_{i}$ from $y_{c}$ and the specific value of $y_{c}$.
Replacing $W_{i}^{r}$ and $W_{i}^{g}$ in Eq. (1) or its hinge loss variant with $W_{i,y_{c}}$ yields the Soft Adaptive Vicinal Discriminator Loss (SAVDL). While SAVDL retains the form of Eq. (1), it fundamentally differs in constructing soft vicinities for training samples.
As illustrated in Fig. 4, this adaptive mechanism dynamically responds to local sample density variations: in dense regions (small $\kappa_{y_{c}}$), weights decay rapidly to maintain label consistency while preserving diversity; in sparse regions (large $\kappa_{y_{c}}$), weights decay slowly to incorporate more distant samples, ensuring training stability and preventing mode collapse.
@@
We further propose a hybrid AV mechanism (Figs. 2(d) and 4) that extends the soft AV approach. This hybrid variant incorporates truncated weight decay by assigning zero weights to samples outside the interval $y_{c}-\kappa_{y_{c}},y_{c}+\kappa_{y_{c}}]$. The hybrid adaptive weight for the $i$-th real/fake sample is computed as:
$\widetilde{W}_{i,y_{c}}=W_{i,y_{c}}\cdot\mathds{1}_{\{|y_{c}-y_{i}|\leq\kappa_{y_{c}}\}}.$ (9)
Similarly, replacing $W_{i}^{r}$ and $W_{i}^{g}$ in Eq. (1) or its hinge loss variant with $\widetilde{W}_{i,y_{c}}$ yields the Hybrid Adaptive Vicinal Discriminator Loss (HAVDL).
By combining the adaptive weighting of soft AV with the boundary constraints of hard vicinity, this approach can further enhances label consistency.
Remark 1 (Algorithm for Adaptive Vicinity Construction).
Result: Vicinity width $\kappa_{y_{c}}$ and weight decay rate $v_{y_{c}}$
1 $\kappa_{l}\leftarrow 0$, $\kappa_{r}\leftarrow 0$;
2 $N_{c}=0$;
3 $t_{l}=i-1$, $t_{r}=i$;
// Vicinity extension
4 while $N_{c}<N_{\text{AV}}$ do
5  if $|y_{(t_{l})}-y_{c}|<|y_{(t_{r})}-y_{c}|$ then
     // Extend leftwards
6     $\kappa_{l}\leftarrow|y_{(t_{l})}-y_{c}|$;
7     $N_{c}\leftarrow N_{c}+N_{t_{l}}$;
8     $t_{l}\leftarrow t_{l}-1$;
9
10  else if $|y_{(t_{l})}-y_{c}|>|y_{(t_{r})}-y_{c}|$ then
     // Extend rightwards
11     $\kappa_{r}\leftarrow|y_{(t_{r})}-y_{c}|$;
12     $N_{c}\leftarrow N_{c}+N_{t_{r}}$;
13     $t_{r}\leftarrow t_{r}+1$;
14
15  else
     // Extend in both directions
16     $\kappa_{l}\leftarrow|y_{(t_{l})}-y_{c}|$;
17     $\kappa_{r}\leftarrow|y_{(t_{r})}-y_{c}|$;
18     $N_{c}\leftarrow N_{c}+N_{t_{l}}+N_{t_{r}}$;
19     $t_{l}\leftarrow t_{l}-1$;
20     $t_{r}\leftarrow t_{r}+1$;
21
22  end if
23
24 end while
[…]
26 $\nu_{y_{c}}\leftarrow 1/\kappa_{y_{c}}$;
Remark 2 (Heuristic Selection of $N_{\text{AV}}$).
The minimum number of effective samples per vicinity ($N_{\text{AV}}$) serves as the crucial hyperparameter in CcGAN-AVAR, with ablation studies on RC-49-I and UTKFace demonstrating optimal performance when $N_{\text{AV}}$ is properly calibrated within a moderate range, as evidenced in Fig. 9.
To determine $N_{\text{AV}}$ in practice, we propose a heuristic approach: First, identify all the distinct training labels
$Y^{u}=\{y_{(1)},...,y_{(M)}\}$ (10)
and their sample counts
$\mathcal{N}=\{N_{1},...,N_{M}\}.$ (11)
$\mathcal{S}=\{S_{1},...,S_{M-1}\}.$ (12)
Finally, compute the averaged sample count
$\bar{S}=\frac{1}{M-1}\sum_{i=1}^{M-1}S_{i},$
and set $N_{\text{AV}}$ to an integer near $\bar{S}$ with potential minor adjustments based on the sparsity characteristics of the data distribution. For instance, on the bimodal RC-49-I dataset where $\bar{S}\approx 25.279$, we set $N_{\text{AV}}=30$. For UTKFace with $\bar{S}\approx 480$, we chose $N_{\text{AV}}=400$. In the case of Steering Angle, where many $S_{i}\in\mathcal{S}$ cluster around 2 despite $\bar{S}\approx 13$, we selected $N_{\text{AV}}=20$ to mitigate the impact of sparse label regions.
III-C Auxiliary Generator Regularization